OpenStudy (anonymous):

Please divide and simplify: x^2-16/x+7 divided by 3x-12/x^2-49

OpenStudy (anonymous):

I would greatly appreciate your help!

OpenStudy (anonymous):

x²-16/x+7 x² - 16 x² - 49 x² - 4² x² - 7² ------------ = ------- * -------- = ------- * -------- 3x-12/x²-49 x + 7 3x - 12 x + 7 3x - 12 (x+4) (x-4) (x+7) (x-7) = ----------- * ----------- x+7 3 (x-4) Cancelling out the common factors from numerator and denominator, we get (x+4) (x-7) x² + 4x - 7x - 28 x² - 3x - 28 = ----------- = ---------------- = ------------- 3 3 3